# Logical perspectives on learning statistical objects

### v1

@@ 8 Related Work
As noted earlier, what we refer to as the“dual distribution class” is from (Hu et al., 2022), while the distribution class is new to our work.
The notions of NIP formulas and structures have been developed over many decades, beginning with work of Shelah, see for example (Shelah, 1990). Many common structures, such as the real field and real exponential field were shown to be NIP (van den Dries, 1998). A fundamental result in (M.C.Laskowski, 1992) is that NIP first order theories are precisely those where definable families are PAC learnable: we use this as a definition of NIP in this work.
The notion of randomization of a structure was introduced in (Keisler, 1999). It was later reformulated in terms of continuous logic by Keisler and Ben Yaacov, see (Ben Yaacov, 2009). (Keisler, 1999; Ben Yaacov & Keisler, 2009) present an extensive study of the randomization transformation, presenting in particular axioms for the randomization. In this paper we deal with a simpler context, and we do not reason across models. Thus our presentation of randomization is simplified.
NIP and other model-theoretic properties were shown to be preserved in moving from a class to its randomization in (Ben Yaacov, 2009; 2013): our bounds refine the analysis given in these works. Characterizations of which partitioned first-order formulas lead to the corresponding concept class being online learnable can be found in (Chase & Freitag, 2019). Model-theoretic characterizations of which formulas are learnable in other learning models (e.g. Private PAC learning) are provided in (Alon et al., 2019).
In the last part of the paper, we deal with prediction and fitting problems for statistical classes: finding parameters that fit training data, or quantifying over these. Such problems have been extensively studied for standard neural architectures (Abrahamsen et al., 2021; Bertschinger et al., 2023; Goel et al., 2021), and also in the presence of SoftMax and other exponential activation functions (Hankala et al., 2023). Our results deal with a different set of function classes, and in the problem we will deal with a fixed hypothesis class, thus fixing the number of parameters that can be set, in contrast to (Abrahamsen et al., 2021; Bertschinger et al., 2023; Goel et al., 2021; Hankala et al., 2023).
@@ 9 Conclusions
We investigated a mapping that takes a “base” hypothesis class, consisting of either Boolean or real-valued functions, to other classes based on probability distributions over either the range space or the parameter space of the class. We connected this to the theory of randomizations in model theory: there we map a classical or continuous-valued structure to another continuous-valued structure based on random variables. We have proved that these transformations preserve agnostic PAC learnability and agnostic online learnability, refining results from both learning of database queries (Hu et al., 2022) and the model theory (Ben Yaacov & Keisler, 2009). In addition to providing a linkage between these communities, our results provide improved bounds. For realizable learning, we have provided counterexamples to preservation. Finally, we provide an initial exploration of the computational complexity of decision problems related to learning distribution classes.
Our motivation concerns distribution classes, but we obtain our positive results by embedding into a more general class, the randomization of a class or structure. This class is strictly more general than the distribution and dual distribution class, since it allows correlation between range elements and parameters. We are currently exploring how to exploit this generality.
@@
Our results on fitting/prediction problems are far from comprehensive, showing only that logical techniques, and more generally model-theoretic properties of the base structure, can be relevant to these problems.
@@ Appendix A Agnostic online learnabilty and duality: Proof of Proposition 1
Although this is surely well known, we include a proof for completeness. This can be done directly using the sequential fat-shattering dimension, Fact 6. However, we can also use the characterization in terms of stability of a CL formula in Theorem 8: although this is stated in the context of a logical formula, any function class can be considered a formula. The theorem states that a function class is online learnable if and only if:
There is some $d$ such that in $\mathfrak{C}$ there are no $a_{1},\dots,a_{d}$, $b_{1},\dots,b_{d}$ such that for all $i<j$,
$|\phi(a_{i};b_{j})-\phi(a_{j};b_{i})|\geq\gamma.$

### v2

@@ 6 Related Work
As noted earlier, what we refer to as the “dual distribution class” is from (Hu et al., 2022), where it is defined only for concept classes. The basic result of (Hu et al., 2022) is that agnostic PAC learnability of a base concept class implies agnostic PAC learnability of the dual distribution class, with accompanying sample complexity bounds that are asymptotically looser than ours. The extension of dual distribution classes to base classes consisting of real-valued values, as well as the notion of the distribution class, is new to our work.
We have presented our arguments without relying on logical notions. But our work is inspired by work on learnability of hypothesis classes coming from logical formulas, so we say more about the connection here. Given a structure $\mathfrak{M}$ and a logical formula $\phi(x_{1}\ldots x_{k})$, one can look at the $k$-tuples from the domain of $\mathfrak{M}$ that satisfy $\phi$: this is a definable set of tuples from $\mathfrak{M}$. Given a first-order formula $\phi(x_{1}\ldots x_{j};y_{1}\ldots y_{k})$ in which the free variables are partitioned, we get a family of sets of $j$-tuples, by varying $\vec{y}$ over all $k$-tuples in $\mathfrak{M}$. Thus each $\phi(\vec{x};\vec{y})$ gives a concept class with range space the $j$-tuples from the domain of $\mathfrak{M}$, and parameter space the $k$-tuples from the structure. For example the family of rectangles in the reals and the family of intervals in the reals are definable families. A partitioned formula is said to be NIP if the corresponding concept class has finite VC dimension.
Model theorists have identified a number of structures where the concept class arising from each partitioned formula has finite VC dimension (or equivalently, is agnostic PAC learnable): these are called NIP structures: see, for example, (Simon, 2015). As discussed in our paper, agnostic online learnability corresponds to a hypothesis class being stable: see Definition 39. A partitioned formula $\phi(\vec{x};\vec{y})$ in a structure $\mathfrak{M}$ is called stable exactly when the corresponding concept class is stable in the sense of Definition 39. The study of stable formulas has been developed over many decades by model theorists, and many structures have been identified in which every partitioned formula is stable: see, for example (Shelah, 1990). There is an analogous notion of “continuous logic”, in which formulas are real-valued, and in continuous logic partitioned formulas give real-valued hypothesis classes. The results in works such as (Ben Yaacov, 2009; Ben Yaacov and Keisler, 2009) are phrased in terms of NIP and stability of continuous logic structures.
The notion of measurable family and its expectation class was introduced in (Ben Yaacov, 2009). It relates to a transformation that is similar to the one we perform here: moving from a logical structure to its “randomization”, another structure where the elements are random variables. The notion of randomization originates in (Keisler, 1999), and was developed further in (Ben Yaacov, 2009; Keisler, 1999; Ben Yaacov and Keisler, 2009). Many of our results on both agnostic PAC learning and agnostic online learning of statistical classes can be seen as refinements of results in these works. Our refinements include: translating these results outside of their original context of hypothesis classes coming from logic to general hypothesis classes, and presenting sample complexity bounds. In the agnostic online learning setting we generalize results proven in prior work only for concept classes to the setting of real-valued hypothesis classes.
Model-theoretic characterizations of which partitioned formulas are learnable in other learning models (e.g. Private PAC learning) are provided in (Alon et al., 2019).
@@ 7 Conclusions
We investigated a mapping that takes a “base” hypothesis class, consisting of either Boolean or real-valued functions, to other classes based on probability distributions over either the range space or the parameter space of the class. We connected this to a theory of randomized families of classes, stemming from work in model theory: there we move from a random hypothesis class to its expectation class. We have proven that these transformations preserve agnostic PAC learnability and agnostic online learnability, refining results from both learning of database queries (Hu et al., 2022) and the model theory (Ben Yaacov and Keisler, 2009). In addition to providing a linkage between these communities, our results provide improved bounds. For realizable learning, we have provided counterexamples to preservation.
Our motivation concerns distribution classes, but we obtain our positive results by embedding into a strictly more general setting, the expectation class of a measurable family. This setting allows correlation between range elements and parameters. We are currently exploring how to exploit this generality.
@@ Appendix A Agnostic online learnabilty and duality: Proof of Proposition 3
Although this is surely well known, we include a proof for completeness. We can use the characterization in terms of stability of a hypothesis class in Theorem 44: The theorem shows that a hypothesis class ${\mathcal{H}}$ on $X$ parameterized by $Y$ is online learnable if and only if:
For all $\gamma>0$, there is some $d$ such that there are no $a_{1},\dots,a_{d}\in X$, $b_{1},\dots,b_{d}\in Y$, such that for all $i<j$,
$|h_{b_{j}}(a_{i})-h_{b_{i}}(a_{j})|\geq\gamma.$